# Dijets with a large rapidity separation at the next-to-leading BFKL for search of large extra dimension gravity at colliders

### v3

@@ VI Results and discussion
Figure 4: The MN dijet cross section with the imposed dijet mass selection $\mbox{$M_{jj}$}>\mbox{$M_{jj\mathrm{min}}$}$ in $pp$ collisions at $\mbox{$\sqrt{s}$}=100$ TeV. The ADD gravity signal is shown for various parameter choices, including the number of extra dimensions $n_{D}$, and the Plank scale $M_{D}$ in the presence of $n_{D}$ extra dimensions. The QCD background includes either contributions calculated with the LO+LL DGLAP or LL/NLL BFKL corrections. Panel (a) corresponds to $\mbox{$M_{jj\mathrm{min}}$}=30$ TeV, and panel (b) to $\mbox{$M_{jj\mathrm{min}}$}=70$ TeV.
@@
The signal calculation for gravity with large extra dimensions, as proposed by the ADD model, is presented for several parameter choices. The number of extra dimensions considered are $n_{D}=2$ and $6$, reflecting typical values studied in this theoretical framework. This allows exploration of how the cross sections and phenomenology depend on the dimensionality of the extra spatial dimensions. The examined values of the Planck scale in the presence of extra dimensions are as follows: $\mbox{$M_{D}$}=1.5$ and $3$ TeV for $\mbox{$\sqrt{s}$}=13$ TeV with $\mbox{$M_{jj}$}>6$ and $9$ TeV (Fig. 2); $\mbox{$M_{D}$}=1.5$ and $3$ TeV for $\mbox{$\sqrt{s}$}=40$ TeV with $\mbox{$M_{jj}$}>9$ TeV (Fig. 3a); $\mbox{$M_{D}$}=10$ and $15$ TeV for $\mbox{$\sqrt{s}$}=40$ TeV with $\mbox{$M_{jj}$}>30$ TeV (Fig. 3b); $\mbox{$M_{D}$}=10$ and $15$ TeV for $\mbox{$\sqrt{s}$}=100$ TeV with $\mbox{$M_{jj}$}>30$ TeV (Fig. 4a); $\mbox{$M_{D}$}=30$ and $40$ TeV for $\mbox{$\sqrt{s}$}=100$ TeV with $\mbox{$M_{jj}$}>70$ TeV (Fig. 4b).
@@
The predictions of the ADD model exhibit less sensitivity to the number of extra dimensions $n_{D}$ and are more strongly impacted by the value of the scale $M_{D}$. The reduced sensitivity of ADD model predictions to the number of extra dimensions $n_{D}$ can be understood as follows. The change of $n_{D}$ at a fixed scale $M_{D}$ alters the size of the extra dimensions, which remain large enough for the particles to be insensitive to this change thanks to the vast difference between particle interaction scales and the scale corresponding to the size of extra dimensions. In contrast, varying $M_{D}$ directly modifies the strength of the gravitational coupling to matter, thus having a more pronounced effect on observable signals.
@@
The QCD background predictions in various approximations are provided with: LO+LL DGLAP approximation; DGLAP-based MC generator pythia8 with the CP5 tune; calculations including LL and NLL BFKL corrections. All of them are presented for all considered collision energies and dijet mass selections. This provides a comprehensive comparison of different theoretical approaches and modeling techniques for the QCD background across a variety of kinematic conditions. The NLL BFKL predictions consistently yield lower estimates for the QCD background compared to other calculations (up to two orders of magnitude). Additionally, the high dijet mass selection $\mbox{$M_{jj}$}>\mbox{$M_{jj\mathrm{min}}$}$ has the most pronounced effect on the NLL BFKL predictions among the various QCD approaches, leading to a stronger suppression of the background in this regime.
@@
Considering the largest rapidity separation currently accessible in experiments at the LHC $\mbox{$\Delta y$}=8.7$ and the background calculated including NLL BFKL corrections, it can be concluded that measurements of MN dijet production with the high dijet mass selection in $pp$ collisions are sensitive to ADD gravity for $\mbox{$M_{D}$}<3$ TeV at $\mbox{$\sqrt{s}$}=13$ TeV; for $\mbox{$M_{D}$}<10$ TeV at $\mbox{$\sqrt{s}$}=40$ TeV; and for $\mbox{$M_{D}$}<30$ TeV at $\mbox{$\sqrt{s}$}=100$ TeV. The sensitivity may improve further if the transition to eikonal regime occurs at lower values of $\Delta y$. The estimated sensitivity of the proposed approach appears to be lower compared to existing searches conducted at the LHC [30, 31, 32, 33, 34, 35, 36]. However, it is important to remember that the LHC searches rely on divergent behavior of effective theory near the scale of new physics. The complementary approach discussed here explores a distinct kinematic regime, thereby it will help to reinforce and validate the conclusions.
To measure a cross section on the order of $\sim 10^{-6}$ pb, an integrated luminosity at least $\sim 1~\mathrm{ab}^{-1}$ is required, which is expected to be available at future facilities such as the HL-LHC, FCC-hh, and CEPC-SppC. However, at high luminosities and large rapidities, it becomes necessary to disentangle overlapping $pp$ collisions occurring within the same or nearby bunch crossings. This imposes stringent requirements for high granularity in the next generation of detectors to effectively resolve and identify large rapidity jets from individual $pp$ collision events and correct for possible overlap effects.
@@ VII Summary
The signal of gravity in the presence of large extra dimensions, as formulated by Arkani-Hamed, Dimopoulos, and Dvali (ADD) [25], is studied for dijet production events with a large rapidity separation within the trans-Planckian eikonal regime defined by $\mbox{$\sqrt{\hat{s}}$}~\gg~\mbox{$M_{D}$}~\gg~\sqrt{-\mbox{$\hat{t}$}}$ in proton-proton ($pp$) collisions. The calculations are performed for experimental conditions projected for future collider facilities such as the HL-LHC, FCCpp and CEPC-SppC, covering $pp$ collision energies ranging from $\mbox{$\sqrt{s}$}=13$ TeV up to $100$ TeV. Parameter values such as the number of extra dimensions $n_{D}=2$ and $6$, and various Planck scales in presence of extra dimensions, $M_{D}$, across collision energies are explored.
@@
The calculations also show that using DGLAP dynamics at large rapidities in the semi-hard regime may significantly overestimate the QCD which could result in misinterpreting experimental data and overlooking potential signals of new physics.

### v4

@@ VI Results and discussion
Figure 4: The MN dijet cross section with the imposed dijet mass selection $\mbox{$M_{jj}$}>\mbox{$M_{jj\mathrm{min}}$}$ in $pp$ collisions at $\mbox{$\sqrt{s}$}=100$ TeV. The ADD gravity signal is shown for various parameter choices, including the number of extra dimensions $n_{D}$, and the Plank scale $M_{D}$ in the presence of $n_{D}$ extra dimensions. The QCD background includes either contributions calculated with the LO+LL DGLAP or LL/NLL BFKL corrections. The systematic uncertainty of the NLL BFKL calculation is denoted by the hatched green band. Panel (a) corresponds to $\mbox{$M_{jj\mathrm{min}}$}=30$ TeV, and panel (b) to $\mbox{$M_{jj\mathrm{min}}$}=70$ TeV.
@@
The signal calculation for gravity with large extra dimensions, as proposed by the ADD model, is presented for several parameter choices. The number of extra dimensions considered are $n_{D}=2$ and $6$, reflecting typical values studied in this theoretical framework. This allows exploration of how the cross sections and phenomenology depend on the dimensionality of the extra spatial dimensions. The examined values of the Planck scale in the presence of extra dimensions are as follows: $\mbox{$M_{D}$}=1.5$ and $3$ TeV for $\mbox{$\sqrt{s}$}=13$ TeV with $\mbox{$M_{jj}$}>6$ and $9$ TeV [Fig. 2]; $\mbox{$M_{D}$}=1.5$ and $3$ TeV for $\mbox{$\sqrt{s}$}=40$ TeV with $\mbox{$M_{jj}$}>9$ TeV [Fig. 3(a)]; $\mbox{$M_{D}$}=10$ and $15$ TeV for $\mbox{$\sqrt{s}$}=40$ TeV with $\mbox{$M_{jj}$}>30$ TeV [Fig. 3(b)]; $\mbox{$M_{D}$}=10$ and $15$ TeV for $\mbox{$\sqrt{s}$}=100$ TeV with $\mbox{$M_{jj}$}>30$ TeV [Fig. 4(a)]; $\mbox{$M_{D}$}=20$ and $30$ TeV for $\mbox{$\sqrt{s}$}=100$ TeV with $\mbox{$M_{jj}$}>70$ TeV [Fig. 4(b)].
@@
The predictions of the ADD model exhibit less sensitivity to the number of extra dimensions $n_{D}$ and are more strongly impacted by the value of the scale $M_{D}$. The reduced sensitivity of ADD model predictions to the number of extra dimensions $n_{D}$ can be understood as follows: the change of $n_{D}$ at a fixed scale $M_{D}$ alters the size of the extra dimensions, which remain large enough for the particles to be insensitive to this change thanks to the vast difference between particle interaction scales and the scale corresponding to the size of extra dimensions. In contrast, varying $M_{D}$ directly modifies the strength of the gravitational coupling to matter, thus having a more pronounced effect on observable signals.
@@
The QCD background predictions in various approximations are provided with: LO+LL DGLAP approximation; DGLAP-based MC generator pythia8 with the CP5 tune; and calculations including LL and NLL BFKL corrections. All of them are presented for all considered collision energies and dijet mass selections. This provides a comprehensive comparison of different theoretical approaches and modeling techniques for the QCD background across a variety of kinematic conditions. The NLL BFKL predictions consistently yield lower estimates for the QCD background compared to other calculations (up to 2 orders of magnitude). Additionally, the high dijet mass selection $\mbox{$M_{jj}$}>\mbox{$M_{jj\mathrm{min}}$}$ has the most pronounced effect on the NLL BFKL predictions among the various QCD approaches, leading to a stronger suppression of the background in this regime.
The systematic uncertainty of NLL BFKL calculation is shown as a hatched green band in Figs. 2, 3, and 4. For the highest $M_{jj\mathrm{min}}$ selections (the most sensitive regions for ADD gravity searches in the trans-Planckian eikonal regime) the total uncertainty is about a factor of  5 (or  1/5) across all collision energies. This remains smaller than the difference between DGLAP and NLL BFKL background estimates.
@@
Considering the largest rapidity separation currently accessible in experiments at the LHC $\mbox{$\Delta y$}=8.7$ and the background calculated including NLL BFKL corrections, it can be concluded that measurements of MN dijet production with the high dijet mass selection in $pp$ collisions are sensitive to ADD gravity for $\mbox{$M_{D}$}<3$ TeV at $\mbox{$\sqrt{s}$}=13$ TeV; for $\mbox{$M_{D}$}<10$ TeV at $\mbox{$\sqrt{s}$}=40$ TeV; and for $\mbox{$M_{D}$}<20$ TeV at $\mbox{$\sqrt{s}$}=100$ TeV. In case the ADD signal is comparable to the QCD background, the large PDF systematic uncertainty can prevent effective separation of the two. More precise PDF measurements at large $x$ are therefore highly desirable. The sensitivity of measurements in trans-Planckin eikonal regime to the ADD gravity may improve further if the transition to eikonal regime occurs at lower values of $\Delta y$.
The estimated sensitivity of the proposed approach appears to be lower compared to existing searches conducted at the LHC [30, 31, 32, 33, 34, 35, 36], which establish a lower bound on the scale of ADD gravity $M_{D}$ at $7\div 10$ TeV. However, it is important to remember that the LHC searches rely on divergent behavior of effective theory near the scale of new physics. The masses of observed dilepton (diphoton) systems reach 3(2.6) TeV which are close to the established lower bound of $M_{D}$. The complementary approach discussed here explores a distinct kinematic regime. In trans-Planckian eikonal regime the scale of gravitons is defined by the $\hat{t}$ variable, which at $\Delta y=8.7$ is $\sim 1\%$ of $M_{jj}$. Thus, the scale of gravitons is about 90 GeV, when $\mbox{$M_{jj}$}=9$ TeV. This scale of gravitons is much lower then considered $M_{D}$ = 3 TeV at 13 TeV collision energy. Studying gravity in the trans-Planckian eikonal regime therefore will help to reinforce and validate the conclusions of existing searches.
An increase in high-mass dijet production at large $\Delta y$ inevitably leads to the simultaneous enhancement of the high-$p_{\perp}$ forward jet spectrum (here, forward denotes large rapidity region). However, calculating single-jet spectra within the BFKL approach remains challenging. Therefore, further development of BFKL calculation methods is needed in search of an ADD gravity signal in forward jet $p_{\perp}$ spectra.
[…]
To measure a cross section on the order of $\sim 10^{-6}$ pb, an integrated luminosity of at least $\sim 1~\mathrm{ab}^{-1}$ is required, which is expected to be available at future facilities such as the HL-LHC, FCC-hh, and CEPC-SppC. However, at high luminosities and large rapidities, it becomes necessary to disentangle overlapping $pp$ collisions occurring within the same or nearby bunch crossings. This imposes stringent requirements for high granularity in the next generation of detectors to effectively resolve and identify large rapidity jets from individual $pp$ collision events and correct for possible overlap effects.
@@ VII Summary
The signal of gravity in the presence of large extra dimensions, as formulated by Arkani-Hamed, Dimopoulos, and Dvali (ADD) [25], is studied for dijet production events with a large rapidity separation within the trans-Planckian eikonal regime defined by $\mbox{$\sqrt{\hat{s}}$}~\gg~\mbox{$M_{D}$}~\gg~\sqrt{-\mbox{$\hat{t}$}}$ in proton-proton collisions. The calculations are performed for experimental conditions projected for future collider facilities such as the HL-LHC, FCCpp, and CEPC-SppC, covering $pp$ collision energies ranging from $\mbox{$\sqrt{s}$}=13$ TeV up to $100$ TeV. Parameter values such as the number of extra dimensions $n_{D}=2$ and $6$, and various Planck scales in presence of extra dimensions, $M_{D}$, across collision energies are explored.
@@
The calculations also show that using DGLAP dynamics at large rapidities in the semi-hard regime may significantly overestimate the QCD which could result in misinterpreting experimental data and overlooking potential signals of new physics. This work demonstrates how the use of NLL BFKL background, rather than a DGLAP extrapolation, can alter the interpretation of dijet measurements with large rapidity separations.
Tabulated results are provided in the Zenodo record [60].
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available [60].